the wave length of Halpha line of Balmer series is X angstram what is the Xof H beta line of Balmer series.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
 
For Hbeta n1 = 2 and n2 = 4 has ebeen defined as well.
So,
1/k = R*1^2*(1/2^2 – 1/4^2) =>R*(1/4 – 1/16) => 3R/16
k = 16/3R would be the required wavelength for sure.